OpenStudy (anonymous):

Find parametric equations for : The line in R^4 that goes through the point (1, 2, 0, 4) and is parallel to the vector (-2, 5, 3, 7).

OpenStudy (anonymous):

So, parallel means that is has the same direction vector right? and through the point (1,2,0,4) just means that you use that as an initial index position valued vector.

OpenStudy (anonymous):

so, is it gonna be x=1-2t y=2+5t z=0+3t and h=4+7t ???

OpenStudy (anonymous):

exactly